# Adapting Pre-Trained Vision Models for Novel Instance Detection and Segmentation

### v2

@@ 4 Experiments
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using InfoNCE loss [67, 15] with temperature 0.05. With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with the Adam optimizer [45] at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout [34] is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the appendix.
4.1 Detection Datasets
@@
High-resolution Dataset. Shen et al. [81] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
@@
According to Shen et al. [81], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [78], RetinaNet [55], CenterNet [104], FCOS [85], and the transformer-based DINO [98]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [11] and DINOv2 [68] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
@@
Synthetic-Real Test Sets. Li et al. [49] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [5] includes 8 texture-less objects and 1,514 bounding boxes, primarily challenging due to significant occlusion. The YCB-Video (YCB-V) [8] features 21 objects and 4,125 bounding boxes, with pose variation as the main challenge. Since these datasets have real testing images without reference videos, Li et al. [49] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
RoboTools Benchmark. Li et al. [49] introduce RoboTools, a challenging real-world benchmark for robotics. It features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. Compared to previous benchmarks [5, 8], RoboTools offers greater difficulty due to its cluttered backgrounds and significant pose variations. Additionally, the benchmark’s reference videos are made from real images and incorporate realistic lighting conditions, including shadows. 25 real template images of each instance are sampled from reference videos.
@@
For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, Li et al. [49] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [42]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [11, 75] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [57]. They also employ the class-level one-shot detectors OS2D [69] and BHRL [95]. To address the limitations of traditional 2D methods with pose variations and occlusions, Li et al. [49] introduced VoxDet which is based on a 3D voxel representation. To train these methods, Li et al. [49] developed the open-world instance detection (OWID-10k) dataset as a synthetic instance detection training set. It features 9,901 objects from ShapeNet [12] and ABO [19], each rendered into a 40-frame, 360-degree video using Blenderproc [22].
@@
Evaluation Metrics. For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05. This metric is the primary evaluation standard in the COCO Object Detection dataset [54]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
4.2 Segmentation Datasets
@@
The BOP Challenge. We test our method using seven datasets from the BOP challenge [84]: LineMod Occlusion (LM-O) [5], T-LESS [35], TUD-L [36], IC-BIN [25], ITODD [26], HomebrewedDB (HB) [40], and YCB-Video [8]. The objects from these datasets vary widely. We use 42 template rendering images from CNOS [66], which are generated via BlenderProc [22].
@@
Figure 4: Visual examples of our results using the Weight Adapter on the high-resolution and RoboTools datasets. Our approach detects specific object instances in cluttered scenes according to their real template images.
Table 2: Detection performance on the fully real dataset, RoboTools. Proposal indicates the object proposal model. GS stands for Grounded-SAM. OLN* is trained with the matching head while OLN employs fixed modules.
@@
Table 3: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. ()($\cdot$)( ⋅ ) represents the number of instances.
@@
Table 4: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances. SAM6D (RGBD) exclusively utilizes RGB-D images, whereas other models employ only RGB images.
@@
Table 5: Comparison of segmentation performance across various proposal models and embedding generation options. The first row is from CNOS [66].
@@
Figure 5: Comparison of segmentation results using CNOS, SAM6D, and our NIDS-Net on the YCB-V dataset. CNOS and SAM6D may misclassify some background regions or object parts as objects due to their reliance on SAM for proposal generation. NIDS-Net addresses this limitation with Grounded-SAM.
4.3 Benchmarking Results
@@
Detection. The high-resolution dataset results are presented in Table 1. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table 2, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
@@
For the Synthetic-Real datasets, LM-O and YCB-V, we present a comparison of our approach with other methods in Table 3. Our model outperforms VoxDet by 10.3 AP on LM-O and by 24.0 AP on YCB-V. The improved performance on YCB-V, which contains more instances, indicates that our adapter functions more effectively with increased instance variety.
@@
Segmentation. CNOS [66] and SAM-6D [53] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table 4. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
@@
Qualitative Result. We present detection results for both simple and complex scenes in Fig. 4. Fig. 5 displays a visual comparison of the segmentation results. Additional examples are available in the appendix.
@@
Runtime. The tables of detection results include the runtime of our method. Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process. Please see the appendix for segmentation runtime.
4.4 Ablation Study
@@
Grounded SAM (GS) vs SAM. On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table 5. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
@@
FFA vs $cls$ token. To compare these two types of embedding generation, we present the results in Table 5. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter. Additional ablation studies are provided in the appendix.
5 Discussions
@@
In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation. We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals. We also present the Weight Adapter, which refines features from a pre-trained vision model effectively and mitigates the risk of overfitting. With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching. Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
@@
However, there are limitations: given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors. Additionally, Grounding DINO may produce large bounding boxes that contain multiple objects, leading to detection failures. Moreover, our method tends to overlook heavily occluded objects with low confidence scores.
@@ Acknowledgments
Supplementary Material
6 Training Details
Detection. For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance. Both adapters are trained for the same number of epochs, as detailed in Table 6. To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
Table 6: The training epochs of different detection datasets.
@@ 6 Training Details
7 More Ablation Study
FFA vs $cls$ token and Grounded SAM (GS) vs SAM in Detection Tasks. The detection capabilities of FFA and the $cls$ token are demonstrated using the high-resolution dataset, as presented in Table 9. For instance embedding generation, FFA achieves significantly greater enhancements compared to the $cls$ token. Furthermore, Table 9 illustrates that GS is effective in generating precise object proposals.
Table 7: Detection results using different ViT backbones of Dinov2. “reg” indicates DINOv2 with registers. The results are based on all testing images of the high-resolution dataset. WA Diff indicates the improvement attributed to the Weight Adapter.
@@ 7 More Ablation Study
Table 8: Detection results on the High-resolution real world detection dataset using various SAM variants.
@@
Table 9: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
@@
Image encoder. Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [81]. As illustrated in Table 10, DINOv2 exhibits superior performance attributable to its robust visual features. Fig. 6 presents the visual results of various image encoders.
Table 10: The instance embeddings of different image encoders.
@@
Dinov2 backbones with adapter. Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table 7. Please refer to the Appendix for additional ablation studies.
Aggregation. For cls token embeddings, averaging the top $k$ highest (avg k) scores yields the best results [66, 53]. For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes. The comparison of these aggregation functions is detailed in Table 11.
SAM variants. We evaluate different SAM variants on the High-resolution real world detection dataset. As illustrated in Table 8, SAM and our weight adapter achieves the superior performance. Given the precise bounding boxes, HQ-SAM[41] yields results comparable to those of SAM [47].
Table 11: Comparison of aggregation functions for segmentation performance. We report Average Precision (AP). avg k refers to averaging the top $k$ scores. All results are based on object proposals from Grounded SAM (GS).
@@
Runtime. We compare the efficiency of existing methods for novel instance segmentation, as presented in Table 12. Our approach significantly reduces running time by proposing only high-quality bounding boxes.
Table 12: Runtime comparisons of various methods for novel instance segmentation.
@@
CNOS FastSAM Tesla V100 0.22
CNOS DeForce RTX 3090 0.23
SAM-6D (RGB) DeForce RTX 3090 0.25
SAM-6D (RGBD) DeForce RTX 3090 0.45
CNOS SAM Tesla V100 1.84
CNOS DeForce RTX 3090 2.35
SAM-6D (RGB) DeForce RTX 3090 2.28
SAM-6D (RGBD) DeForce RTX 3090 2.80
NIDS-Net (Ours) GS RTX A5000 0.49
@@
8 Unseen Detection of BOP Benchmark
We compare our approach with ZeroPose [13], CNOS [66], and SAM-6D [53] for 2D unseen detection, as illustrated in Table 13. Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
Table 13: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).

### v3

@@ IV Experiments
The experiments are conducted on an RTX A5000 GPU. For each dataset, we train both adapters on template embeddings using the InfoNCE loss [11, 12]. With two linear layers, both adapters initially decrease the embedding dimension from $C$ to $C/4$, subsequently restoring it to $C$. The Weight Adapter is trained with the Adam optimizer [31] at a learning rate of 1e-3 and a batch size of 1024, while the CLIP-Adapter is trained at a learning rate of 1e-4. Dropout [32] is incorporated in the CLIP-Adapter to reduce overfitting. For additional details, please see the appendix.
IV-A Detection Datasets
@@
High-resolution Dataset. [2] design a high-resolution dataset for instance detection. This dataset comprises 100 distinct object instances, each represented by 24 photos taken from multiple viewpoints, with each photo having a resolution of $3072\times 3072$ pixels. These instances are integrated into 14 different indoor scenes for testing, captured in even higher resolution ($6144\times 8192$ pixels). The testing images are further classified as easy or hard based on the level of scene clutter and how much the instances are occluded. Easy tags are assigned when objects are sparsely placed, while hard tags are used for more cluttered setups.
@@
According to [2], two types of baselines are set for the high-resolution dataset: Cut-Paste-Learn and a non-learned method. Cut-Paste-Learn generates synthetic training images with 2D-box annotations, putting foreground instances in various sizes and aspect ratios onto different backgrounds. This enables training detectors by viewing each instance as a class. They train five detectors: FasterRCNN [25], RetinaNet [26], CenterNet [27], FCOS [28], and the transformer-based DINO [29]. For the non-learned method, SAM is initially used to generate proposals, followed by employing DINO [30] and DINOv2 [7] to generate features for both proposal and template images, ultimately performing proposal matching and selection.
@@
Synthetic-Real Test Sets. [1] employ two benchmarks for evaluation. LineMod-Occlusion (LM-O) [33] includes 8 texture-less objects and 1,514 bounding boxes. The YCB-Video (YCB-V) [34] features 21 objects and 4,125 bounding boxes. Since these datasets have real testing images without reference videos, [1] generate synthetic videos using CAD models. We sample 16 synthetic template images per object from these videos.
@@
RoboTools Benchmark [1] features 20 distinct instances, 9,109 annotations, and 24 complex scenarios. 25 real template images per instance are sampled from their reference videos. For cluttered scenes from the datasets RoboTools, LM-O and YCB-V, [1] have developed several 2D baselines: OLNDINO, OLNCLIP, and OLNCorr. Initially, they generate open-world 2D proposals using their detection module [35]. For matching, different methods are employed to select the proposal with the highest score. In OLNDINO and OLNCLIP, they use robust features from pre-trained backbones [30, 36] and cosine similarity for matching. For OLNCorr, a matching head is designed based on correlation [20]. They also employ the class-level one-shot detectors OS2D [37] and BHRL [38]. To address the limitations of traditional 2D methods with pose variations and occlusions, [1] introduced VoxDet which is based on a 3D voxel representation. To train these methods, [1] developed a synthetic instance detection dataset (OWID-10k).
@@
Evaluation Metrics. For detection, we evaluate our method with Average Precision (AP). AP is computed by averaging precision scores at various Intersection over Union (IoU) thresholds, specifically from 0.5 to 0.95, in increments of 0.05 [39]. Additionally, AP50 and AP75 are variations of this metric, where precision is averaged across all instances at IoU thresholds of 0.5 and 0.75, respectively.
IV-B Segmentation Datasets
@@
The BOP Challenge. We test our method using seven datasets from the BOP challenge [15]: LineMod Occlusion (LM-O) [33], T-LESS [40], TUD-L [41], IC-BIN [42], ITODD [43], HomebrewedDB (HB) [44], and YCB-Video [34]. We use 42 template rendering images from CNOS [3], which are generated via BlenderProc [45].
@@
TABLE II: Detection performance on the fully real dataset, RoboTools. Proposal indicates the object proposal model. GS stands for Grounded-SAM. OLN* is trained with the matching head while OLN employs fixed modules.
@@
TABLE III: Detection performance on the LM-O and YCB-V datasets. OLN* is trained alongside the matching head whereas OLN uses fixed modules. †means the model is trained on OWID and real images. ()($\cdot$)( ⋅ ) represents the number of instances.
@@
TABLE IV: Novel instance segmentation results on the seven core datasets of the BOP benchmark. We utilize Average Precision (AP) to compare these methods. ()($\cdot$)( ⋅ ) includes the number of instances. SAM6D (RGBD) exclusively utilizes RGB-D images, whereas other models employ only RGB images.
@@
TABLE V: Segmentation performance comparison across proposal models and embedding methods. The first row is from CNOS [3].
@@
Figure 5: Comparison of segmentation results using CNOS, SAM6D, and NIDS-Net on the YCB-V dataset. CNOS and SAM6D may misclassify some background regions or object parts as objects due to proposal generation of SAM. Red arrows indicate these mistakes. NIDS-Net addresses this limitation with Grounded-SAM.
IV-C Benchmarking Results
@@
Detection. The high-resolution dataset results are presented in Table I. Our models dramatically outperform existing techniques, achieving the highest AP scores across all categories. Our basic method surpasses the top baseline by 17.7 AP. Additionally, as this dataset contains 100 instances with some similarities among them, our Weight Adapter boosts overall performance by 4.6 AP and improves the detection of small objects by 5.4 AP. For the RoboTools dataset, detailed in Table II, our method outperforms the state-of-the-art VoxDet by 46.2 AP. Since RoboTools contains only 20 instances, our adapter has limited scope to enhance performance. Our approach achieves over 60 AP on these two fully real datasets.
@@
For the Synthetic-Real datasets, LM-O and YCB-V, we present their results in Table III. Our model outperforms VoxDet by 10.3 AP on LM-O and by 24.0 AP on YCB-V. The improved performance on YCB-V, which contains more instances, indicates that our adapter functions more effectively with increased instance variety.
@@
Segmentation. CNOS [3] and SAM-6D [4] initially utilize SAM or FastSAM to obtain proposals and derive instance embeddings through the $cls$ token of Dinov2. In contrast, we generate proposals using Grounded-SAM and acquire embeddings through the FFA pipeline. The segmentation results for the seven principal datasets of the BOP challenge are detailed in Table IV. Overall, our method surpasses the state-of-the-art SAM-6D and achieves superior results compared to RGBD results.
@@
Qualitative Results. We present detection results of RoboTools in Fig. 4. Moreover, Fig. 5 displays a visual comparison of the segmentation results. Additional examples are available in the appendix.
@@
Runtime. Table I includes the runtime of our method for object detection. Grounded-SAM substantially decreases the number of proposals compared to SAM, thereby accelerating the entire process. Please see the appendix for segmentation runtime.
IV-D Real-world Testing and Failure Cases
We capture 14 multi-view template images per object using seven Intel RealSense D455 cameras for real-world objects. With a trained weight adapter on these images, our method is tested on these objects with one D455 camera, demonstrating robust performance across various scenes (Fig. 6). However, it misclassifies objects when highly similar instances are present as in Fig. 7. With the weight adapter trained on BOP datasets, we also tested our method on YCB-V objects using a Fetch robot. The demonstrations are included in the supplementary video.
Figure 6: Evaluation of NIDS-Net on real-world objects across diverse scenes. All predictions are accurate in these examples.
Figure 7: Examples of failure cases from NIDS-Net, where the orange arrows indicate incorrect predictions.
TABLE VI: The performance comparison of $\beta$ in the Weight Adapter on the high-resolution dataset.
$\beta$ 1 5 10 20 50 100
AP 62.0 63.1 63.9 63.1 62.7 62.2
AP50 73.7 75.6 76.6 75.6 75.1 74.3
AP75 68.3 69.5 70.6 69.3 69.0 68.2
IV-E Ablation Study
@@
Grounded-SAM (GS) vs SAM. On the seven BOP datasets containing images of numerous cluttered scenes, we compare the object proposals from GS and SAM as detailed in Table V. The results indicate that GS yields more precise bounding boxes and masks compared to using SAM alone. Furthermore, GS enhances efficiency by eliminating false object proposals, thereby reducing runtime.
@@
FFA vs $cls$ token. To compare these two types of embedding generation, we present the results in Table V. Our weight adapter enhances both two types of embeddings. Despite close segmentation results, FFA produces embeddings that possess greater adaptive potential, demonstrated by higher AP scores after adaption via our weight adapter.
$\beta$ in our Weight Adapter. We evaluate various values of $\beta$ on the high-resolution dataset. All results presented in Table VI surpass those obtained using CLIP-Adapter. We set $\beta$ to to 10 across all datasets as it yields the best results. Additional ablation studies are provided in the appendix.
V Discussions
@@
In this study, we introduce NIDS-Net, a framework designed for novel instance detection and segmentation. We utilize Grounding DINO and SAM to generate precise foreground object proposals, as opposed to conventional naive region proposals. We also introduce the Weight Adapter, which effectively refines features from a pre-trained vision model, mitigates the risk of overfitting, and adjusts the weighting of cosine similarities. With the adapter, template and proposal embeddings of different instances are separated to facilitate the subsequent matching. Our method surpasses other approaches significantly in detection performance and also excels in segmentation compared to existing methods.
@@
However, there are limitations. Given that our approach incorporates multiple pre-trained models, it requires greater computational resources compared to end-to-end detectors. When instances exhibit highly similar appearances, NIDS-Net may encounter detection failures. The method sometimes missed heavily occluded objects with low confidence scores.
@@ ACKNOWLEDGMENT
APPENDIX
VI Training Details
Detection. For detection datasets, the weight adapter is trained with a batch size of 1024, while the CLIP-Adapter is trained with a batch size of 512 to enhance performance. Both adapters are trained for the same number of epochs, as detailed in Table VII. To utilize Grounding DINO, a box threshold of 0.15 is set for the high-resolution and YCB-V datasets, and 0.10 for other datasets.
TABLE VII: The training epochs of different detection datasets.
@@ VI Training Details
VII More Ablation Study
TABLE VIII: Detection results using different ViT backbones of Dinov2. “reg” indicates DINOv2 with registers. The results are based on all testing images of the high-resolution dataset. WA Diff indicates the improvement attributed to the Weight Adapter.
@@ VII More Ablation Study
TABLE IX: Detection results on the High-resolution real-world detection dataset using various SAM variants.
@@
TABLE X: The detection results on all images of the high-resolution dataset. “Proposal” refers to the object proposal method. “Embedding” denotes the method of instance embedding generation.
@@
Image encoder. Given the same object proposals with GS, we evaluate the FFA embeddings from different image encoders on the High-resolution dataset [2]. As illustrated in Table XI, DINOv2 exhibits superior performance attributable to its robust visual features. Fig. 8 presents the visual results of various image encoders.
TABLE XI: The instance embeddings of different image encoders.
@@
Dinov2 backbones with adapter. Our Weight Adapter is compatible with various backbones of Dinov2. Notably, more powerful backbones, which offer a more effective feature space, enable our adapter to deliver greater improvements. Details of this comparison are provided in Table VIII.
Aggregation. For cls token embeddings, averaging the top $k$ highest (avg k) scores yields the best results [3, 4]. For FFA embeddings, the $\max$ aggregation function achieves optimal outcomes. The comparison of these aggregation functions is detailed in Table XII.
SAM variants. We evaluate different SAM variants on the High-resolution real world detection dataset. As illustrated in Table IX, SAM and our weight adapter achieves the superior performance. Given the precise bounding boxes, HQ-SAM[49] yields results comparable to those of SAM [5].
TABLE XII: Comparison of aggregation functions for segmentation performance. We report Average Precision (AP). avg k refers to averaging the top $k$ scores. All results are based on object proposals from Grounded SAM (GS).
@@
Runtime. We compare the efficiency of existing methods for novel instance segmentation, as presented in Table XIII. Our approach significantly reduces running time by proposing only high-quality bounding boxes.
TABLE XIII: Runtime comparisons of various methods for novel instance segmentation.
@@
CNOS [3] FastSAM Tesla V100 0.22
CNOS [3] DeForce RTX 3090 0.23
SAM-6D (RGB) [4] DeForce RTX 3090 0.25
SAM-6D (RGBD) [4] DeForce RTX 3090 0.45
CNOS [3] SAM Tesla V100 1.84
CNOS [3] DeForce RTX 3090 2.35
SAM-6D (RGB) [4] DeForce RTX 3090 2.28
SAM-6D (RGBD) [4] DeForce RTX 3090 2.80
NIDS-Net w/o adapter (Ours) GS RTX A5000 0.49
@@
VIII Unseen Detection of BOP Benchmark
We compare our approach with ZeroPose [46], CNOS [3], and SAM-6D [4] for 2D unseen detection, as illustrated in Table XIV. Our method outperforms the best RGB method by 2.5 AP and competes effectively with the top RGB-D method.
TABLE XIV: Unseen instance detection results across the seven core datasets of the BOP benchmark, with all results reported as Average Precision (AP).